# Diffusion Bridge Implicit Models

### v1

@@ 4 Sampling with Generalized Diffusion Bridges
Now that we have confirmed the rationality and built the theoretical foundations for applying the generalized diffusion bridge $p_{\theta}$ (Eqn. (14)) to pretrained DDBMs, a range of inference processes is now at our disposal, controlled by the variance parameter $\rho$. This positions us to explore the resultant sampling procedures and the effects of $\rho$ in pursuit of better and more efficient generation.
@@ 4.1 Diffusion Bridge Implicit Models
Suppose we sample in reverse time on the discretized timesteps $0=t_{0}<t_{1}<\dots<t_{N-1}<t_{N}=T$. The number $N$ and the schedule of sampling steps can be made independently of the original timesteps on which the bridge model is trained, whether discrete [25] or continuous [53]. According to the generative process of $p_{\theta}$ in Eqn. (13), the updating rule from $t_{n+1}$ to $t_{n}$ is described by
@@
Intuitively, the form of Eqn. (17) bears a resemblance to the forward transition kernel of the diffusion bridge in Eqn. (6) (which can be rewritten as $\bm{x}_{t}=a_{t}\bm{x}_{T}+b_{t}\bm{x}_{0}+c_{t}\bm{\epsilon},\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$). Compared to the forward process, $\bm{x}_{0}$ is substituted with the predicted $\hat{\bm{x}}_{0}$, and a portion of the standard Gaussian noise $\bm{\epsilon}$ now stems from the predicted noise $\hat{\bm{\epsilon}}$. The predicted noise $\hat{\bm{\epsilon}}$ is derived from $\bm{x}_{t_{n+1}}$ at the previous timestep and can be expressed by the predicted clean data $\hat{\bm{x}}_{0}$.
@@
DDPM [16] ScoreSDE [44] DDIM [40] DDBM [53] DBIM (Ours)
Timesteps Discrete Continuous Discrete Continuous Discrete
Forward Distribution $q(\bm{x}_{n}|\bm{x}_{0})$ $q(\bm{x}_{t}|\bm{x}_{0})$ $q(\bm{x}_{n}|\bm{x}_{n-1},\bm{x}_{0})$ $q(\bm{x}_{t}|\bm{x}_{0},\bm{x}_{T})$ $q(\bm{x}_{t_{n+1}}|\bm{x}_{0},\bm{x}_{t_{n}},\bm{x}_{T})$
Inference Process $p_{\theta}(\bm{x}_{n-1}|\bm{x}_{n})$ SDE/ODE $p_{\theta}(\bm{x}_{n-1}|\bm{x}_{n})$ SDE/ODE $p_{\theta}(\bm{x}_{t_{n}}|\bm{x}_{t_{n+1}},\bm{x}_{T})$
Markovian
We investigate the effects of the variance parameter $\rho$ from the theoretical perspective by considering two extreme cases. Firstly, we note that when $\rho_{n}=\sigma_{t_{n}}\sqrt{1-\frac{\text{SNR}_{t_{n+1}}}{\text{SNR}_{t_{n}}}}$ for each $10\leq n\leq N-$ 1, the $\bm{x}_{T}$ term in Eqn. (17) is canceled out. In this scenario, the forward process in Eqn. (4.1) becomes a Markovian bridge (see details in Appendix C). Besides, the inference process will get rid of $\bm{x}_{T}$ and simplify to $p_{\theta}(\bm{x}_{t_{n}}|\bm{x}_{t_{n+1}})$, akin to the sampling mechanism in DDPMs [16]. Secondly, when $\rho_{n}=0$ for each $10\leq n\leq N-$ 1, the inference process will be free from random noise and composed of deterministic iterative updates, characteristic of an implicit probabilistic model [31]. Consequently, we name the resulting model diffusion bridge implicit models (DBIMs), drawing parallels with denoising diffusion implicit models (DDIMs) [40], a widely adopted technique to expedite the sampling of diffusion models. DBIMs serve as the bridge counterpart and extension of DDIMs, as illustrated in Table 1.
@@
To tackle this, we are forced to employ the other boundary choice $\rho_{n}=\sigma_{t_{n}}\sqrt{1-\frac{\text{SNR}_{t_{n+1}}}{\text{SNR}_{t_{n}}}}$ in the initial step [^2] With this choice, $c_{t_{n+1}}$ in the denominator in Eqn. (17) will be canceled out., which is aligned with our previous restriction that $\rho_{N-1}=c_{t_{N-1}}$. This will introduce an additional standard Gaussian noise $\bm{\epsilon}$ which we term as the booting noise. It accounts for the stochasticity of the final sample $\bm{x}_{0}$ under a given fixed $\bm{x}_{T}$ and can be viewed as the latent variable. We illustrate the complete DBIM pipeline in Figure 2.
@@ 4.2 Connection to Probability Flow ODE
It is intuitive to perceive that the deterministic sampling can be related to solving an ODE. By setting $\rho=0,t_{n+1}=t$ and $t_{n+1}-t_{n}=\Delta t$ in Eqn. (17), the DBIM updating rule can be reorganized as
$\frac{\bm{x}_{t-\Delta t}}{c_{t-\Delta t}}=\frac{\bm{x}_{t}}{c_{t}}+\left(\frac{a_{t-\Delta t}}{c_{t-\Delta t}}-\frac{a_{t}}{c_{t}}\right)\bm{x}_{T}+\left(\frac{b_{t-\Delta t}}{c_{t-\Delta t}}-\frac{b_{t}}{c_{t}}\right)\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$ (18)
As $a_{t},b_{t},c_{t}$ are continuous functions of time $t$ defined in Eqn. (6), the ratios $\frac{a_{t}}{c_{t}}$ and $\frac{b_{t}}{c_{t}}$ also remain continuous functions of $t$. Therefore, Eqn. (18) can be treated as an Euler discretization of the following ordinary differential equation (ODE):
@@
We retain this form due to its extreme neatness. Though it does not resemble a conventional ODE involving $\mathrm{d}t$, the two infinitesimal terms $\mathrm{d}\left(\frac{a_{t}}{c_{t}}\right)$ and $\mathrm{d}\left(\frac{b_{t}}{c_{t}}\right)$ can indeed be expressed with $\mathrm{d}t$ by the chain rule of derivatives. The ODE form of DBIMs also suggests that with a sufficient number of discretization steps, we can reverse the sampling process and obtain encodings of the observed data, which can be useful for interpolation or other downstream tasks.
In DDBMs, a probability flow ODE (PF-ODE, Eqn. (8)) involving $\mathrm{d}\bm{x}_{t}$ and $\mathrm{d}t$ is proposed and used for deterministic sampling. We reveal in the following proposition that our ODE in Eqn. (19) can exactly yield the PF-ODE, while the derivation of our ODE does not rely on any advanced knowledge of differential equations (such as the Fokker-Planck equation), but pure probabilistic models.
@@
Suppose $\bm{s}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$ is learned as the ground-truth bridge score $\nabla_{\bm{x}_{t}}\log q(\bm{x}_{t}|\bm{x}_{T})$, and $\bm{x}_{\theta}$ is related to $\bm{s}_{\theta}$ through Eqn. (16), then Eqn. (19) can be converted to the probability flow ODE (Eqn. (8)) proposed in DDBMs.
Though the conversion from our ODE to the PF-ODE is straightforward, the reverse conversion can be non-trivial and may require utilizing more advanced tools such as exponential integrators [4, 18]. Despite their theoretical equivalence, our ODE describes the evolution of $\frac{\bm{x}_{t}}{c_{t}}$ rather than $\bm{x}_{t}$, and its discretization is performed with respect to $\mathrm{d}\left(\frac{a_{t}}{c_{t}}\right)$ and $\mathrm{d}\left(\frac{b_{t}}{c_{t}}\right)$ instead of $\mathrm{d}t$. Empirically, DBIMs prove significantly more efficient than the Euler discretization of the PF-ODE, thereby accelerating DDBMs by a substantial margin.
@@ 5 Experiments
where $\eta=0$ corresponds to deterministic sampling and $\eta=1$ corresponds to Markovian stochastic sampling. Note that while in diffusion models, the Markovian case of DDIMs degrades to DDPMs (i.e., the original sampler for discrete-time diffusion models), DDBMs are initially defined over continuous time and sampled by simulating differential equations. Consequently, DBIMs with $\eta=1$ also serve as a novel sampling scheme for DDBMs.
@@ 5.1 Sample Quality and Efficiency
Pix2Pix [19] 1 74.8 3.24 0.356 0.209 82.4 4.22 0.556 0.133
DDIB [45] $\geq 40$ 186.84 2.04 0.869 1.05 242.3 4.22 0.798 0.794
SDEdit [30] $\geq 40$ 26.5 3.58 0.271 0.510 31.14 5.70 0.714 0.534
Rectified Flow [26] $\geq 40$ 25.3 2.80 0.241 0.088 77.18 5.87 0.534 0.157
I2SB [25] $\geq 40$ 7.43 3.40 0.244 0.191 9.34 5.77 0.373 0.145
DDBM [53] 118 1.83 3.73 0.142 0.040 4.43 6.21 0.244 0.084
DDBM [53] 200 0.88 3.69 0.110 0.006 3.34 5.95 0.215 0.020
@@
DDRM [22] 20 24.4 62.1
$\Pi$GDM [41] 100 7.3 72.6
DDNM [48] 100 15.1 55.9
Palette [35] 1000 6.1 63.0
I2SB [25] 1000 4.9 66.1
DDBM [53] 500 4.27 71.8
@@
The comparison of visual quality is illustrated in Figure 1 and Figure 3, where DBIM produces smoother outputs with significantly fewer noisy artifacts compared to DDBM’s hybrid sampler. Additional samples are provided in Appendix E.
@@
Ablation of the Variance Parameter We investigate the impact of the variance parameter $\rho$ (controlled by $\eta$) to identify how the level of stochasticity affects sample quality across various NFEs, as shown in Table 4 and Table 5. For image translation tasks, we consistently observe that employing a deterministic sampler with $\eta=0.0$ yields superior performance compared to stochastic samplers with $\eta>0$. We attribute it to the characteristics of the datasets, where the target image is highly correlated with and dependent on the condition, resulting in a generative model that lacks diversity. In this case, a straightforward mapping without the involvement of stochasticity is preferred. Conversely, for image inpainting on the more diverse dataset ImageNet 256$\times$256, the parameter $\eta$ exhibits significance across different NFEs. When $\text{NFE}\leq 20$, $\eta=0.0$ is the optimal choice, with FID steadily increasing as $\eta$ ascends. However, when $\text{NFE}\geq 50$, a relatively large level of stochasticity at $\eta=0.8$ or even $\eta=1.0$ yields optimal FID. Notably, the FID of $\eta=0.0$ converges to 3.95 at $\text{NFE}=100$, with no further improvement at larger NFEs, indicating convergence to the ground-truth sample by the corresponding PF-ODE. This observation aligns with diffusion models, where deterministic sampling facilitates rapid convergence, while introducing stochasticity in sampling enhances diversity, ultimately culminating in the highest sample quality when NFE is substantial.
Figure 4: Encoding and reconstruction with deterministic DBIMs.
Figure 5: Semantic interpolation with deterministic DBIMs.

### v2

@@ 4 Sampling with Generalized Diffusion Bridges
Now that we have confirmed the rationality and built the theoretical foundations for applying the generalized diffusion bridge $p_{\theta}$ to pretrained DDBMs, a range of inference processes is now at our disposal, controlled by the variance parameter $\rho$. This positions us to explore the resultant sampling procedures and the effects of $\rho$ in pursuit of better and more efficient generation.
@@ 4.1 Diffusion Bridge Implicit Models
Suppose we sample in reverse time on the discretized timesteps $0=t_{0}<t_{1}<\dots<t_{N-1}<t_{N}=T$. The number $N$ and the schedule of sampling steps can be made independently of the original timesteps on which the bridge model is trained, whether discrete (Liu et al., 2023) or continuous (Zhou et al., 2023). According to the generative process of $p_{\theta}$ in Eqn. (12), the updating rule from $t_{n+1}$ to $t_{n}$ is described by
@@
Intuitively, the form of Eqn. (15) resembles the forward transition kernel of the diffusion bridge in Eqn. (6) (which can be rewritten as $\bm{x}_{t}=a_{t}\bm{x}_{T}+b_{t}\bm{x}_{0}+c_{t}\bm{\epsilon},\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$). In comparison, $\bm{x}_{0}$ is substituted with the predicted $\hat{\bm{x}}_{0}$, and a portion of the standard Gaussian noise $\bm{\epsilon}$ now stems from the predicted noise $\hat{\bm{\epsilon}}$. The predicted noise $\hat{\bm{\epsilon}}$ is derived from $\bm{x}_{t_{n+1}}$ at the previous timestep and can be expressed by the predicted clean data $\hat{\bm{x}}_{0}$.
@@
DDPM (Ho et al., 2020) ScoreSDE (Song et al., 2021c) DDIM (Song et al., 2021a) I2SB (Liu et al., 2023) DDBM (Zhou et al., 2023) DBIM (Ours)
Noise Schedule VP Any Any VE Any Any
Timesteps Discrete Continuous Discrete Discrete Continuous Discrete
Forward Distribution $q(\bm{x}_{n}|\bm{x}_{0})$ $q(\bm{x}_{t}|\bm{x}_{0})$ $q(\bm{x}_{n}|\bm{x}_{n-1},\bm{x}_{0})$ $q(\bm{x}_{n}|\bm{x}_{0},\bm{x}_{N})$ $q(\bm{x}_{t}|\bm{x}_{0},\bm{x}_{T})$ $q(\bm{x}_{t_{n+1}}|\bm{x}_{0},\bm{x}_{t_{n}},\bm{x}_{T})$
Inference Process $p_{\theta}(\bm{x}_{n-1}|\bm{x}_{n})$ SDE/ODE $p_{\theta}(\bm{x}_{n-1}|\bm{x}_{n})$ $p_{\theta}(\bm{x}_{n-1}|\bm{x}_{n})$ SDE/ODE $p_{\theta}(\bm{x}_{t_{n}}|\bm{x}_{t_{n+1}},\bm{x}_{T})$
Non-Markovian
Effects of the Variance Parameter
We investigate the effects of the variance parameter $\rho$ from the theoretical perspective by considering two extreme cases. Firstly, we note that when $\rho_{n}=\sigma_{t_{n}}\sqrt{1-\frac{\text{SNR}_{t_{n+1}}}{\text{SNR}_{t_{n}}}}$ for each $10\leq n\leq N-$ 1, the $\bm{x}_{T}$ term in Eqn. (15) is canceled out. In this scenario, the forward process in Eqn. (4.1) becomes a Markovian bridge (see details in Appendix C.1). Besides, the inference process will get rid of $\bm{x}_{T}$ and simplify to $p_{\theta}(\bm{x}_{t_{n}}|\bm{x}_{t_{n+1}})$, akin to the sampling mechanism in DDPMs (Ho et al., 2020). Secondly, when $\rho_{n}=0$ for each $10\leq n\leq N-$ 1, the inference process will be free from random noise and composed of deterministic iterative updates, characteristic of an implicit probabilistic model (Mohamed & Lakshminarayanan, 2016). Consequently, we name the resulting model diffusion bridge implicit models (DBIMs), drawing parallels with denoising diffusion implicit models (DDIMs) (Song et al., 2021a). DBIMs serve as the bridge counterpart and extension of DDIMs, as illustrated in Table 1.
@@
To be theoretically robust, we employ the other boundary choice $\rho_{n}=\sigma_{t_{n}}\sqrt{1-\frac{\text{SNR}_{t_{n+1}}}{\text{SNR}_{t_{n}}}}$ in the initial step [^2] With this choice, $c_{t_{n+1}}$ in the denominator in Eqn. (15) will be canceled out., which is aligned with our previous restriction that $\rho_{N-1}=c_{t_{N-1}}$. This will introduce an additional standard Gaussian noise $\bm{\epsilon}$ which we term as the booting noise. It accounts for the stochasticity of the final sample $\bm{x}_{0}$ under a given fixed $\bm{x}_{T}$ and can be viewed as the latent variable. We illustrate the complete DBIM pipeline in Figure 2.
@@ 4.2 Connection to Probability Flow ODE
It is intuitive to perceive that the deterministic sampling can be related to solving an ODE. By setting $\rho=0,t_{n+1}=t$ and $t_{n+1}-t_{n}=\Delta t$ in Eqn. (15), the DBIM updating rule can be reorganized as $\frac{\bm{x}_{t-\Delta t}}{c_{t-\Delta t}}=\frac{\bm{x}_{t}}{c_{t}}+\left(\frac{a_{t-\Delta t}}{c_{t-\Delta t}}-\frac{a_{t}}{c_{t}}\right)\bm{x}_{T}+\left(\frac{b_{t-\Delta t}}{c_{t-\Delta t}}-\frac{b_{t}}{c_{t}}\right)\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$. As $a_{t},b_{t},c_{t}$ are continuous functions of time $t$ defined in Eqn. (6), the ratios $\frac{a_{t}}{c_{t}}$ and $\frac{b_{t}}{c_{t}}$ also remain continuous functions of $t$. Therefore, DBIM ($\rho=0$) can be treated as an Euler discretization of the following ordinary differential equation (ODE):
@@
Though it does not resemble a conventional ODE involving $\mathrm{d}t$, the two infinitesimal terms $\mathrm{d}\left(\frac{a_{t}}{c_{t}}\right)$ and $\mathrm{d}\left(\frac{b_{t}}{c_{t}}\right)$ can be expressed with $\mathrm{d}t$ by the chain rule of derivatives. The ODE form also suggests that with a sufficient number of discretization steps, we can reverse the sampling process and obtain encodings of the observed data, which can be useful for interpolation or other downstream tasks.
In DDBMs, the PF-ODE (Eqn. (8)) involving $\mathrm{d}\bm{x}_{t}$ and $\mathrm{d}t$ is proposed and used for deterministic sampling. We reveal in the following proposition that our ODE in Eqn. (16) can exactly yield the PF-ODE without relying on the advanced Kolmogorov forward (or Fokker-Planck) equation.
@@
Suppose $\bm{s}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$ is learned as the ground-truth bridge score $\nabla_{\bm{x}_{t}}\log q(\bm{x}_{t}|\bm{x}_{T})$, and $\bm{x}_{\theta}$ is related to $\bm{s}_{\theta}$ through Eqn. (14), then Eqn. (16) can be converted to the PF-ODE (Eqn. (8)) proposed in DDBMs.
Though the conversion from our ODE to the PF-ODE is straightforward, the reverse conversion can be non-trivial and require complex tools such as exponential integrators (Calvo & Palencia, 2006; Hochbruck et al., 2009) (Appendix C.4). We highlight our differences from the PF-ODE in DDBMs: (1) Our ODE has a novel form with exceptional neatness. (2) Despite their theoretical equivalence, our ODE describes the evolution of $\frac{\bm{x}_{t}}{c_{t}}$ rather than $\bm{x}_{t}$, and its discretization is performed with respect to $\mathrm{d}\left(\frac{a_{t}}{c_{t}}\right)$ and $\mathrm{d}\left(\frac{b_{t}}{c_{t}}\right)$ instead of $\mathrm{d}t$. (3) Empirically, DBIMs ($\rho=0$) prove significantly more efficient than the Euler discretization of the PF-ODE, thereby accelerating DDBMs by a substantial margin. (4) In contrast to the fully deterministic ODE, DBIMs are capable of various stochastic levels to achieve the best generation quality under the same sampling steps.
We discuss some special cases of DBIM and their connection to related works in Appendix C.3.
4.3 Extension to High-Order Methods
The simplicity and efficiency of our ODE (Eqn. (16)) also inspire novel high-order numerical solvers tailored for DDBMs, potentially bringing faster convergence than the first-order Euler discretization. Specifically, using the time change-of-variable $\lambda_{t}=\log\left(\frac{b_{t}}{c_{t}}\right)=\frac{1}{2}\left(\text{SNR}_{t}-\text{SNR}_{T}\right)$, the solution of Eqn. (16) from time $t$ to time $s<t$ can be represented as
$\bm{x}_{s}=\frac{c_{s}}{c_{t}}\bm{x}_{t}+\left(a_{s}-\frac{c_{s}}{c_{t}}a_{t}\right)\bm{x}_{T}+c_{s}\int_{\lambda_{t}}^{\lambda_{s}}e^{\lambda}\bm{x}_{\theta}(\bm{x}_{t_{\lambda}},t_{\lambda},\bm{x}_{T})\mathrm{d}\lambda$ (17)
where $t_{\lambda}$ is the inverse function of $\lambda_{t}$. The intractable integral can be approximated by Taylor expansion of $\bm{x}_{\theta}$ and finite difference estimations of high-order derivatives, following well-established numerical methods (Hochbruck & Ostermann, 2005) and their extensive application in diffusion models (Zhang & Chen, 2022; Lu et al., 2022b; Gonzalez et al., 2023). We present the derivations of our high-order solvers in Appendix D, and the detailed algorithm in Appendix E.
@@ 5 Experiments
where $\eta=0$ and $\eta=1$ correspond to deterministic sampling and Markovian stochastic sampling.
@@ 5.1 Sample Quality and Efficiency
Pix2Pix (Isola et al., 2017) 1 74.8 3.24 0.356 0.209 82.4 4.22 0.556 0.133
DDIB (Su et al., 2022) $\geq 40$ 186.84 2.04 0.869 1.05 242.3 4.22 0.798 0.794
SDEdit (Meng et al., 2022) $\geq 40$ 26.5 3.58 0.271 0.510 31.14 5.70 0.714 0.534
Rectified Flow (Liu et al., 2022) $\geq 40$ 25.3 2.80 0.241 0.088 77.18 5.87 0.534 0.157
I2SB (Liu et al., 2023) $\geq 40$ 7.43 3.40 0.244 0.191 9.34 5.77 0.373 0.145
DDBM (Zhou et al., 2023) 118 1.83 3.73 0.142 0.040 4.43 6.21 0.244 0.084
DDBM (Zhou et al., 2023) 200 0.88 3.69 0.110 0.006 3.34 5.95 0.215 0.020
@@
DDRM (Kawar et al., 2022) 20 24.4 62.1
$\Pi$GDM (Song et al., 2023a) 100 7.3 72.6
DDNM (Wang et al., 2023) 100 15.1 55.9
Palette (Saharia et al., 2022) 1000 6.1 63.0
I2SB (Liu et al., 2023) 1000 4.9 66.1
DDBM (Zhou et al., 2023) 500 4.27 71.8
@@
The comparison of visual quality is illustrated in Figure 1 and Figure 3, where DBIM produces smoother outputs with significantly fewer noisy artifacts compared to DDBM’s hybrid sampler. Additional samples are provided in Appendix G.
@@
Ablation of the Variance Parameter We investigate the impact of the variance parameter $\rho$ (controlled by $\eta$) to identify how the level of stochasticity affects sample quality across various NFEs, as shown in Table 4 and Table 5. For image translation tasks, we consistently observe that employing a deterministic sampler with $\eta=0$ yields superior performance compared to stochastic samplers with $\eta>0$. We attribute it to the characteristics of the datasets, where the target image is highly correlated with and dependent on the condition, resulting in a generative model that lacks diversity. In this case, a straightforward mapping without the involvement of stochasticity is preferred. Conversely, for image inpainting on the more diverse dataset ImageNet 256$\times$256, the parameter $\eta$ exhibits significance across different NFEs. When $\text{NFE}\leq 20$, $\eta=0$ is the optimal choice, with FID steadily increasing as $\eta$ ascends. However, when $\text{NFE}\geq 50$, a relatively large level of stochasticity at $\eta=0.8$ or even $\eta=1$ yields optimal FID. Notably, the FID of $\eta=0$ converges to 3.95 at $\text{NFE}=100$, with no further improvement at larger NFEs, indicating convergence to the ground-truth sample by the corresponding PF-ODE. This observation aligns with diffusion models, where deterministic sampling facilitates rapid convergence, while introducing stochasticity in sampling enhances diversity, ultimately culminating in the highest sample quality when NFE is substantial.
Table 6: The effects of high-order methods, measured by FID.
Sampler NFE
5 10 20 50 100 5 10 20 50 100 5 10 20 50 100
Image Translation Inpainting
[…]
DBIM ($\eta=0$) 3.60 2.46 1.74 1.14 0.89 14.25 7.98 4.99 3.20 2.57 6.47 4.62 4.17 4.00 3.95
DBIM (2nd-order) 3.44 2.15 1.46 0.97 0.78 13.70 7.23 4.37 2.87 2.41 5.85 4.45 4.15 4.00 3.97
DBIM (3rd-order) 3.38 2.09 1.43 0.95 0.77 13.42 7.03 4.22 2.85 2.40 5.83 4.42 4.12 4.00 3.96
High-Order Methods
We further demonstrate the effects of high-order methods by comparing them to deterministic DBIM, the first-order case. As shown in Table 6, high-order methods consistently improve FID scores in image translation tasks, as well as in inpainting tasks when NFE$<50$, resulting in enhanced generation quality in the low NFE regime. Besides, the 3rd-order variant performs slightly better than the 2nd-order variant. However, in contrast to the numerical solvers in diffusion models, the benefits of high-order extensions are relatively minor in diffusion bridges and less pronounced than the improvement when adjusting $\eta$ from 1 to 0. Nevertheless, high-order DBIMs are significantly more efficient than DDBM’s PF-ODE-based high-order solvers.
As illustrated in Figure 1, our high-order sampler produces images of similar semantic content to the first-order case, using the same booting noise. In contrast, the visual quality is improved with finer textures, resulting in better FID. This indicates that the high-order gradient information from past network outputs benefits the generation quality by adding high-frequency visual details.